# Deciding Local Unitary Equivalence of Graph States in Quasi-Polynomial Time

### v1

@@ 3.3 An algorithm to recognise LCr-equivalent graph states
A multiset $S$ of vertices of type X can be represented as a vector in $({\mathbb{Z}/2^{r}\mathbb{Z}})^{V_{X}}$ which entries are $S(u)\bmod 2^{r}$, as the multiplicity can be considered modulo $2^{r}$ in the context of a $r$-local complementation. Let $\Sigma$ be the subset of $({\mathbb{Z}/2^{r}\mathbb{Z}})^{V_{X}}$ which corresponds to all $r$-incident independent multisets of vertices in $V_{X}$. $\Sigma$ is a vector space since the property of $r$-incidence is preserved under the addition of two vectors. $\Sigma$ is actually the space of solutions to a set of $O(n^{r+1})$ equations [^6] There are precisely $\binom{|V_{X}|}{r+1}+\binom{|V_{X}|}{r}+\cdots+\binom{|V_{X}|}{3}+\binom{|V_{X}|}{2}$ equations. given by the conditions of $r$-incidence. With some multiplications by powers of 2, all these equations are expressible as equations modulo $2^{r}$. For every set $K\subseteq V\setminus\textup{{supp}}(S)$ of size between $2$ and $r+1$, the corresponding equation is
@@
Let $f$ be the function that associates with each element $S\in\Sigma$, its action $f(S)\in\Omega$ on the edges: for any $u,v\in V_{Z}$, $f(S)_{u,v}=u\sim_{G_{1}}v\leavevmode\nobreak\ \oplus\leavevmode\nobreak\ u\sim_{G_{1}\star^{r}S}v$. Notice that $f$ is linear, as for any $S,S^{\prime}\in\Sigma$:
$f(S+S^{\prime})$ $=u\sim_{G_{1}}v\leavevmode\nobreak\ \oplus\leavevmode\nobreak\ u\sim_{G_{1}\star^{r}(S+S^{\prime})}v=u\sim_{G_{1}}v\leavevmode\nobreak\ \oplus\leavevmode\nobreak\ u\sim_{(G_{1}\star^{r}S)\star^{r}S^{\prime}}v$
$=u\sim_{G_{1}}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ u\sim_{G_{1}\star^{r}S}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ \left(S^{\prime}\bullet\Lambda_{G_{1}\star^{r}S}^{u,v}=2^{r-1}\bmod 2^{r}\right)$
$=u\sim_{G_{1}}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ u\sim_{G_{1}\star^{r}S}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ \left(S^{\prime}\bullet\Lambda_{G_{1}}^{u,v}=2^{r-1}\bmod 2^{r}\right)$
$=u\sim_{G_{1}\star^{r}S}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ u\sim_{G_{1}\star^{r}S^{\prime}}v$
$=u\sim_{G_{1}}v\leavevmode\nobreak\ \oplus\leavevmode\nobreak\ u\sim_{G_{1}\star^{r}S}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ u\sim_{G_{1}}v\leavevmode\nobreak\ \oplus\leavevmode\nobreak\ u\sim_{G_{1}\star^{r}S^{\prime}}v=f(S)+f(S^{\prime})$
This directly implies that $\Omega$ is a vector space: if $\omega,\omega^{\prime}\in\Omega$, by definition there exists $S,S^{\prime}\in\Sigma$ such that $f(S)=\omega$ and $f(S^{\prime})=\omega$, moreover $\omega+\omega^{\prime}=f(S)+f(S^{\prime})=f(S+S^{\prime})\in\Omega$. Also, let us prove that $\Omega$ is generated by $\{f(S_{1}),f(S_{2}),...,f(S_{t})\}$. Take a vector $\tilde{\omega}\in\Omega$, by definition there exists $\tilde{S}\in\Sigma$ such that $\tilde{\omega}=f(\tilde{S})$. $\tilde{S}$ can be expressed as a linear combination of vectors from the generating set, i.e. $\tilde{S}=\sum_{i\in[1,t]}a_{i}S_{i}$ where $a_{i}\in\mathbb{Z}/2^{r}\mathbb{Z}$. Then, for any $u,v\in V_{Z}$, $(\tilde{\omega})_{u,v}=f\left(\sum_{i\in[1,t]}a_{i}S_{i}\right)_{u,v}=\sum_{i\in[1,t]}a^{\prime}_{i}f(S_{i})_{u,v}$ by linearity of $f$, where $a^{\prime}_{i}\in\mathbb{F}_{2}$ such that $a^{\prime}_{i}=a_{i}\bmod 2$ when $a_{i}$ and $a^{\prime}_{i}$ are viewed as integers. In other words, $\tilde{w}=\sum_{i\in[1,t]}a^{\prime}_{i}f(S_{i})$, implying that $\{f(S_{1}),f(S_{2}),...,f(S_{t})\}$ is a generating set of $\Omega$. Using Gaussian elimination, one can easily obtain a basis $\mathcal{B}$ of $\Omega$ from $\{f(S_{1}),f(S_{2}),...,f(S_{t})\}$.
Thanks to the exhaustive description of all possible $r$-local complementations on $G_{1}$, we are now ready to reduce LCr-equivalence to LC-equivalence with some additional constraint. We denote by $G_{1}^{\#}$ (resp. $G_{2}^{\#}$) the graph obtained from $G_{1}$ (resp. $G_{2}$) by the following procedure. First, remove the vertices of $V_{X}$. Then, for each vector $\omega\in\mathcal{B}$, for each $u,v\in V_{Z}$ such that $\omega_{u,v}=1$, add a vertex connected only to $u$ and $v$ and call it $p_{u,v}^{\omega}$, and let $\mathcal{P}^{\omega}=\{p_{u,v}^{\omega}\leavevmode\nobreak\ |\leavevmode\nobreak\ \omega_{u,v}=1\}$. In the following, we refer to the vertices added by this procedure as "new vertices".
@@
The proof of Lemma 3.17 makes use of Lemma 3.14 and is given in Appendix D.
@@
Deciding whether there exists a sequence of (possibly repeating) vertices $a_{1},\cdots,a_{m}$ such that $G^{\#}_{2}=G^{\#}_{1}\star a_{1}\star\cdots\star a_{m}$, satisfying the additional constraints described in Lemma 3.17, can be done in runtime $O(n^{4})$.
@@
If $\Omega$ is not of dimension zero, then $G^{\#}_{1}$ and $G^{\#}_{2}$ both have at least one even-degree vertex (since every "new vertex" is of degree 2). Using the notations of Section 3.1, let us define the following linear constraints on $V^{4}$: $\forall u\in V_{Z}$, $\forall\omega\in\mathcal{B}$, $\forall v,v^{\prime}\in\mathcal{P}^{\omega}$,
@@
According to Proposition 3.5, a solution to the system of equations composed of (i), (ii) (see Proposition 3.1) and the additional linear constraints can be computed in runtime $O(n^{4})$ when it exists. Following the algorithm in Appendix A, such a solution yields a sequence of local complementations satisfying the additional constraints described in Lemma 3.17. Conversely, such a sequence of local complementations can be converted into a valid solution to the system of equations.
@@
    Decide whether ${G}^{\#}_{1}$ and ${G}^{\#}_{2}$ are LC-equivalent with the additional constraints described in Lemma 3.17. Output YES if this is the case, NO otherwise.
@@ 3.4 Bounds for generalised local complementation
[…]
@@
By contradiction assume $S$ is an $r$-incident independent multiset which is not genuine. Let $S^{\prime}$ be the multiset obtained from $S$ by choosing, for every set $K\subseteq V\setminus\textup{{supp}}(S)$ s.t. $\{u\in\textup{{supp}}(S)\leavevmode\nobreak\ |\leavevmode\nobreak\ N_{G}(u)=K\}$ is not empty, a single vertex $u\in\textup{{supp}}(S)$ s.t. $N_{G}(u)=K$, and setting $S^{\prime}(u)=\sum_{N_{G}(u)=K}S(u)$ and for any other vertex $v\in\textup{{supp}}(S)$ s.t. $N_{G}(v)=K$, $S^{\prime}(v)=0$. It is direct to show that $S^{\prime}$ is $r$-incident and that $G\star^{r}S=G\star^{r}S^{\prime}$. Then, let $S^{\prime}/2$ be the multiset obtained from $S^{\prime}$ by dividing by 2 the multiplicity of each vertex in $\textup{{supp}}(S^{\prime})$. It is direct to show that $S^{\prime}$ is $(r-1)$-incident and that $G\star^{r}S=G\star^{r-1}S^{\prime}/2$.
@@
$=\sum_{N_{G^{\prime}}(u)=K_{1}}S(u)+|\{K\subseteq K_{0}\leavevmode\nobreak\ |\leavevmode\nobreak\ K_{1}\subsetneq K\}|\bmod 2\text{\leavevmode\nobreak\ \leavevmode\nobreak\ by hypothesis of induction}$
@@
Thus, $\sum_{N_{G^{\prime}}(u)=K_{1}}S(u)$ is odd. As a consequence, for any $K\subseteq K_{0}$ s.t. $|K|>1$, there exists at least one vertex $u\in\textup{{supp}}(S)$ s.t. $N_{G^{\prime}}(u)=K$. Then, $|\textup{{supp}}(S)|\geqslant|\{K\subseteq K_{0}\leavevmode\nobreak\ |\leavevmode\nobreak\ |K|>1\}|=2^{m}-m-1\geqslant 2^{r+2}-(r+2)-1=2^{r+2}-r-3$.
Likewise, $r$-local complementations that cannot be implemented by $(r-1)$-local complementations can only occur or multisets with sufficiently many vertices outside their support.
@@
The proof of Lemma 3.26 involves similar techniques as the proof of Lemma 3.24 and is given in Appendix E. Lemmas 3.24 and 3.26 together give a simple bound involving only the order of the graph.
@@
Put differently, any $r$-local complementation on a graph of order at most $12^{r+2}-$ 1 can be implemented by $(r-1)$-local complementations:
@@
If two graphs of order at most $12^{r+2}-$ 1 are LCr-equivalent, then they are LCr-1-equivalent.
In other words, two LCr-equivalent but not LCr-1-equivalent graphs are of order at least $2^{r+2}$. This implies the following strengthening of Theorem 2.5.
@@
If two graphs of order at most $12^{r+3}-$ 1 are LU-equivalent, they are LCr-equivalent.
@@
Suppose that $G_{1}$ and $G_{2}$ of order $n\leqslant 2^{r+3}-1$ are LU-equivalent. According to Theorem 2.5, $G_{1}$ and $G_{2}$ are LC⌊n/2⌋-1-equivalent. If $\lfloor n/2\rfloor-1\leqslant r$ then $G_{1}$ and $G_{2}$ are trivially LCr-equivalent. Otherwise, according to Corollary 3.28, $G_{1}$ and $G_{2}$ are LCr-equivalent by direct induction.
@@ 3.5 An algorithm to recognise LU-equivalent graph states
According to Theorem 3.20, we have an algorithm that recognises two LCr-equivalent graphs of order $n$ in runtime $O(rn^{r+2.38}+n^{6,38})$. According to Corollary 3.29, $G_{1}$ and $G_{2}$ are LU-equivalent if and only if they are LCr-equivalent, where $r=\log_{2}(n)+O(1)$. Thus, our algorithm that decides LCr-equivalence translates directly to an algorithm that decides LU-equivalence.
@@
In comparison, Burchardt et al. algorithm for LU-equivalence [6] has two sources of exponential time complexity. The logarithmic upper bound on the level of generalised local complementation we introduce may mitage one of these sources (making one parameter of the complexity quasi-polynomial), but does not affect a priori the second one, which is roughly speaking the number of connected components of an intersection graph related to the MLS cover.
@@ 4 LU- and LC-equivalence coincide for graph states up to 19 qubits
The best bound known so far [^9] In [6] it is proved that the number of LU- and LC-orbits of unlabelled graphs of order up to 11 is the same. is that for graphs of order up to 8, LU=LC i.e. LU- and LC-equivalence coincide [7]. The results of Section 3.4 (see Corollary 3.29) already imply a substantial improvement on this bound: LU=LC for graphs of order up to 15. Furthermore, for graphs of order up to 31, LU=LC2, i.e. if two graphs of order up to 31 are LU-equivalent, they are LC2-equivalent. Thus, asking whether LU=LC holds for graphs of order up to 26 is equivalent to asking whether LC2=LC holds for graphs of order up to 26. One direction is to study when a 2-local complementation on an multiset $S$ can be implemented using only usual local complementations over vertices in the support of $S$. If this were to be the case for every graph of order up to 26, it would show that the 27-vertex counterexample is minimal in number of vertices. In the following we study the structure of 2-local complementation to prove that LU=LC holds for graph of order up to 19.
According to Lemma 3.26, if there are at most 4 vertices outside the support of some 2-incident independent multiset $S$, then a 2-local complementation on $S$ can be implemented by usual local complementations. In the peculiar case of 2-local complementation, we are able to use computer-assisted generation (see Appendix F) to extend the result. The code is available at [9].
@@
Likewise, according to Lemma 3.24 and Proposition 3.22, if the support of some 2-incident independent multiset $S$ is of size at most 10, then a 2-local complementation on $S$ can be implemented by usual local complementations. To extend this result to 2-incident independent multisets whose supports is of size at most 12, we first study the case of twin-less sets (two distinct non-connected vertices $u$ and $v$ are twins if $N_{G}(u)=N_{G}(v)$).
@@
The proof of Lemma 4.2 is an induction over the number of vertices connected to $S$ and is given in Appendix G.
According to Proposition 2.3, any 2-local complementation can be decomposed into 1- and 2-local complementations over sets. Furthermore, one can check that if an 2-incident independent set $S$ contains two twins $u$ and $v$, then a 2-local complementation over $S$ has the same effect as a 2-local complementation over $S\setminus\{u,v\}$ followed by a local complementation over $u$. Thus, the action of a 2-local complementation can be described by a 2-local complementation over a twin-less set followed by usual local complementations. Then, Lemma 4.2 can be applied on the twin-less set to yield the following result:
@@
According to Lemma 4.1 and Lemma 4.3, if a 2-incident independent multiset $S$ satisfies $|supp(S)|\leqslant 12$ or $|V\setminus\textup{{supp}}(S)|\leqslant 5$, or alternatively if $|supp(S)|\leqslant 20$ and $|V\setminus\textup{{supp}}(S)|=6$, then a 2-local complementation over $S$ can be implemented by usual local complementations. Thus, for graphs of order up to 19, any 2-local complementation can be implemented by usual local complementations, implying LU=LC. In other words, a 2-local complementation that cannot be implemented by usual local complementation is possible only on a graph of order at least 20. We summarise our findings in the following proposition:

### v3

@@ 3.3 An algorithm to recognise LCr-equivalent graph states
A multiset $S$ of vertices of type X can be represented as a vector in $({\mathbb{Z}/2^{r}\mathbb{Z}})^{V_{X}}$ which entries are $S(u)\bmod 2^{r}$, as the multiplicity can be considered modulo $2^{r}$ in the context of a $r$-local complementation. Let $\Sigma$ be the subset of $({\mathbb{Z}/2^{r}\mathbb{Z}})^{V_{X}}$ that corresponds to all $r$-incident independent multisets of vertices in $V_{X}$. $\Sigma$ is a vector space since the property of $r$-incidence is preserved under the addition of two vectors. $\Sigma$ is actually the space of solutions to a set of $O(n^{r+1})$ equations [^6] There are precisely $\binom{|V_{Z}|}{r+1}+\binom{|V_{Z}|}{r}+\cdots+\binom{|V_{Z}|}{3}+\binom{|V_{Z}|}{2}$ equations. given by the conditions of $r$-incidence. With some multiplications by powers of 2, all these equations are expressible as equations modulo $2^{r}$. For every set $K\subseteq V_{Z}$ of size between $2$ and $r+1$, the corresponding equation is
@@
Let $f$ be the function that associates with each element $S\in\Sigma$, its action $f(S)\in\Omega$ on the edges: for any $u,v\in V_{Z}$, $f(S)_{u,v}=u\sim_{G_{1}}v\penalty 10000\ \oplus\penalty 10000\ u\sim_{G_{1}\star^{r}S}v$. Notice that $f$ is linear, as for any $S,S^{\prime}\in\Sigma$:
$f(S+S^{\prime})$ $=u\sim_{G_{1}}v\penalty 10000\ \oplus\penalty 10000\ u\sim_{G_{1}\star^{r}(S+S^{\prime})}v=u\sim_{G_{1}}v\penalty 10000\ \oplus\penalty 10000\ u\sim_{(G_{1}\star^{r}S)\star^{r}S^{\prime}}v$
$=u\sim_{G_{1}}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ u\sim_{G_{1}\star^{r}S}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ \left(S^{\prime}\bullet\Lambda_{G_{1}\star^{r}S}^{u,v}=2^{r-1}\bmod 2^{r}\right)$
$=u\sim_{G_{1}}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ u\sim_{G_{1}\star^{r}S}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ \left(S^{\prime}\bullet\Lambda_{G_{1}}^{u,v}=2^{r-1}\bmod 2^{r}\right)$
$=u\sim_{G_{1}\star^{r}S}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ u\sim_{G_{1}\star^{r}S^{\prime}}v$
$=u\sim_{G_{1}}v\penalty 10000\ \oplus\penalty 10000\ u\sim_{G_{1}\star^{r}S}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ u\sim_{G_{1}}v\penalty 10000\ \oplus\penalty 10000\ u\sim_{G_{1}\star^{r}S^{\prime}}v=f(S)+f(S^{\prime})$
[…]
Thanks to the exhaustive description of all possible $r$-local complementations on $G_{1}$, we are now ready to reduce LCr-equivalence to LC-equivalence with some additional constraint. We denote by $G_{1}^{\#}$ (resp. $G_{2}^{\#}$) the graph obtained from $G_{1}$ (resp. $G_{2}$) by the following procedure. First, remove the vertices of $V_{X}$. Then, for each vector $\omega\in\mathcal{B}$, for each $u,v\in V_{Z}$ such that $\omega_{u,v}=1$, add a vertex connected only to $u$ and $v$ and call it $p_{u,v}^{\omega}$, and let $\mathcal{P}^{\omega}=\{p_{u,v}^{\omega}\penalty 10000\ |\penalty 10000\ \omega_{u,v}=1\}$. In the following, we refer to the vertices added by this procedure as "new vertices".
@@
The proof of Lemma˜3.17 makes use of Lemma˜3.14 and is given in Appendix˜D.
@@
Deciding whether there exists a sequence of (possibly repeating) vertices $a_{1},\cdots,a_{m}$ such that $G^{\#}_{2}=G^{\#}_{1}\star a_{1}\star\cdots\star a_{m}$, satisfying the additional constraints described in Lemma˜3.17, can be done in runtime $O(n^{4})$.
@@
If $\Omega$ is not of dimension zero, then $G^{\#}_{1}$ and $G^{\#}_{2}$ both have at least one even-degree vertex (since every "new vertex" is of degree 2). Using the notations of Section˜3.1, let us define the following linear constraints on $V^{4}$: $\forall u\in V_{Z}$, $\forall\omega\in\mathcal{B}$, $\forall v,v^{\prime}\in\mathcal{P}^{\omega}$,
@@
According to Proposition˜3.5, a solution to the system of equations composed of (i), (ii) (see Proposition˜3.1) and the additional linear constraints can be computed in runtime $O(n^{4})$ when it exists. Following the algorithm in Appendix˜A, such a solution yields a sequence of local complementations satisfying the additional constraints described in Lemma˜3.17. Conversely, such a sequence of local complementations can be converted into a valid solution to the system of equations.
@@
    Decide whether ${G}^{\#}_{1}$ and ${G}^{\#}_{2}$ are LC-equivalent with the additional constraints described in Lemma˜3.17. Output YES if this is the case, NO otherwise.
@@ 3.4 Bounds for generalised local complementation
In this section, we prove an upper bound on the level of a valid generalised local complementation: roughly speaking we show that if $G\star^{r}S$ is valid then $r$ is at most logarithmic in the order of the graph $G$. This bound is however not true in general as it has been shown in [11] that whenever $G\star^{r}S$ is valid, we have $G\star^{r}S=G\star^{r+1}(S+S)$. To avoid these pathological cases, we thus focus on genuine $r$-incident independent multisets:
@@
By contradiction assume $S$ is an $r$-incident independent multiset that is not genuine. Let $S^{\prime}$ be the multiset obtained from $S$ by choosing, for every set $K\subseteq V\setminus\textup{{supp}}(S)$ s.t. $\{u\in\textup{{supp}}(S)\penalty 10000\ |\penalty 10000\ N_{G}(u)=K\}$ is not empty, a single vertex $u\in\textup{{supp}}(S)$ s.t. $N_{G}(u)=K$, and setting $S^{\prime}(u)=\sum_{N_{G}(u)=K}S(u)$ and for any other vertex $v\in\textup{{supp}}(S)$ s.t. $N_{G}(v)=K$, $S^{\prime}(v)=0$. It is direct to show that $S^{\prime}$ is $r$-incident and that $G\star^{r}S=G\star^{r}S^{\prime}$. Then, let $S^{\prime}/2$ be the multiset obtained from $S^{\prime}$ by dividing by 2 the multiplicity of each vertex in $\textup{{supp}}(S^{\prime})$. It is direct to show that $S^{\prime}$ is (r1)(r-1)-incident and that $G\star^{r}S=G\star^{r-1}S^{\prime}/2$.
@@
$=\sum_{N_{G^{\prime}}(u)=K_{1}}S(u)+|\{K\subseteq K_{0}\penalty 10000\ |\penalty 10000\ K_{1}\subsetneq K\}|\bmod 2\text{\penalty 10000\ \penalty 10000\ by hypothesis of induction}$
@@
Thus, $\sum_{N_{G^{\prime}}(u)=K_{1}}S(u)$ is odd. As a consequence, for any $K\subseteq K_{0}$ s.t. $|K|>1$, there exists at least one vertex $u\in\textup{{supp}}(S)$ s.t. $N_{G^{\prime}}(u)=K$. Then, $|\textup{{supp}}(S)|\geqslant|\{K\subseteq K_{0}\penalty 10000\ |\penalty 10000\ |K|>1\}|=2^{m}-m-1\geqslant 2^{r+2}-(r+2)-1=2^{r+2}-r-3$.
Likewise, $r$-local complementations that cannot be implemented by (r1)(r-1)-local complementations can only occur or multisets with sufficiently many vertices outside their support.
@@
The proof of Lemma˜3.26 involves similar techniques as the proof of Lemma˜3.24 and is given in Appendix˜E. Lemmas˜3.24 and 3.26 together give a simple bound involving only the order of the graph.
@@
Put differently, any $r$-local complementation on a graph of order at most $2^{r+2}-1$ can be implemented by (r1)(r-1)-local complementations:
@@
If two graphs of order at most $2^{r+2}-1$ are LCr-equivalent, then they are LCr-1-equivalent.
In other words, two LCr-equivalent but not LCr-1-equivalent graphs are of order at least $2^{r+2}$. This implies the following strengthening of Theorem˜2.6.
@@
If two graphs of order at most $2^{r+3}-1$ are LU-equivalent, they are LCr-equivalent.
@@
Suppose that $G_{1}$ and $G_{2}$ of order $n\leqslant 2^{r+3}-1$ are LU-equivalent. According to Theorem˜2.6, $G_{1}$ and $G_{2}$ are LC⌊n/2⌋-1-equivalent. If $\lfloor n/2\rfloor-1\leqslant r$ then $G_{1}$ and $G_{2}$ are trivially LCr-equivalent. Otherwise, according to Corollary˜3.28, $G_{1}$ and $G_{2}$ are LCr-equivalent by direct induction.
@@ 3.5 An algorithm to recognise LU-equivalent graph states
According to Theorem˜3.20, we have an algorithm that recognises two LCr-equivalent graphs of order $n$ in runtime $O(rn^{r+2.38}+n^{6,38})$. According to Corollary˜3.29, $G_{1}$ and $G_{2}$ are LU-equivalent if and only if they are LCr-equivalent, where $r=\log_{2}(n)+O(1)$. Thus, our algorithm that decides LCr-equivalence translates directly to an algorithm that decides LU-equivalence.
@@
In comparison, Burchardt et al. algorithm for LU-equivalence [6] has two sources of exponential time complexity. The logarithmic upper bound on the level of generalised local complementation we introduce may mitigate one of these sources (making one parameter of the complexity quasi-polynomial), but does not affect a priori the second one, which is roughly speaking the number of connected components of an intersection graph related to the MLS cover.
@@ 4 LU- and LC-equivalence coincide for graph states up to 19 qubits
The best bound known so far [^8] In [6] it is proved that the number of LU- and LC-orbits of unlabelled graphs of order up to 11 is the same. is that for graphs of order up to 8, LU=LC i.e. LU- and LC-equivalence coincide [7]. The results of Section˜3.4 (see Corollary˜3.29) already imply a substantial improvement on this bound: LU=LC for graphs of order up to 15. Furthermore, for graphs of order up to 31, LU=LC2, i.e. if two graphs of order up to 31 are LU-equivalent, they are LC2-equivalent. Thus, asking whether LU=LC holds for graphs of order up to 26 is equivalent to asking whether LC2=LC holds for graphs of order up to 26. One direction is to study when a 2-local complementation on an multiset $S$ can be implemented using only usual local complementations over vertices in the support of $S$. If this were to be the case for every graph of order up to 26, it would show that the 27-vertex counterexample is minimal in number of vertices. In the following we study the structure of 2-local complementation to prove that LU=LC holds for graph of order up to 19.
According to Lemma˜3.26, if there are at most 4 vertices outside the support of some 2-incident independent multiset $S$, then a 2-local complementation on $S$ can be implemented by usual local complementations. In the peculiar case of 2-local complementation, we are able to use computer-assisted generation (see Appendix˜F) to extend the result. The code is available at [9].
@@
Likewise, according to Lemma˜3.24 and Proposition˜3.22, if the support of some 2-incident independent multiset $S$ is of size at most 10, then a 2-local complementation on $S$ can be implemented by usual local complementations. To extend this result to 2-incident independent multisets whose supports is of size at most 12, we first study the case of twin-less sets (two distinct non-connected vertices $u$ and $v$ are twins if $N_{G}(u)=N_{G}(v)$).
@@
The proof of Lemma˜4.2 is an induction over the number of vertices connected to $S$ and is given in Appendix˜G.
According to Proposition˜2.4, any 2-local complementation can be decomposed into 1- and 2-local complementations over sets. Furthermore, one can check that if an 2-incident independent set $S$ contains two twins $u$ and $v$, then a 2-local complementation over $S$ has the same effect as a 2-local complementation over $S\setminus\{u,v\}$ followed by a local complementation over $u$. Thus, the action of a 2-local complementation can be described by a 2-local complementation over a twin-less set followed by usual local complementations. Then, Lemma˜4.2 can be applied on the twin-less set to yield the following result:
@@
According to Lemma˜4.1 and Lemma˜4.3, if a 2-incident independent multiset $S$ satisfies $|supp(S)|\leqslant 12$ or $|V\setminus\textup{{supp}}(S)|\leqslant 5$, or alternatively if $|supp(S)|\leqslant 20$ and $|V\setminus\textup{{supp}}(S)|=6$, then a 2-local complementation over $S$ can be implemented by usual local complementations. Thus, for graphs of order up to 19, any 2-local complementation can be implemented by usual local complementations, implying LU=LC. In other words, a 2-local complementation that cannot be implemented by usual local complementation is possible only on a graph of order at least 20. We summarise our findings in the following proposition: